# Identifying Important Group of Pixels using Interactions

### v1

@@ Abstract
To better understand the behavior of image classifiers, it is useful to visualize the contribution of individual pixels to the model prediction. In this study, we propose a method, MoXI (Model eXplanation by Interactions), that efficiently and accurately identifies a group of pixels with high prediction confidence. The proposed method employs game-theoretic concepts, Shapley values and interactions, taking into account the effects of individual pixels and the cooperative influence of pixels on model confidence. Theoretical analysis and experiments demonstrate that our method better identifies the pixels that are highly contributing to the model outputs than widely-used visualization methods using Grad-CAM, Attention rollout, and Shapley value. While prior studies have suffered from the exponential computational cost in the computation of Shapley value and interactions, we show that this can be reduced to linear cost for our task.
@@ 1 Introduction
Visualization of important image pixels has been widely used to understand machine learning models in computer vision tasks such as image classification [31, 23, 1, 20, 18, 3]. To this end, visualization methods compute the contribution of each pixel to model decisions.
@@
For example, Grad-CAM [23] measures the contribution using a weighted sum of the feature maps of convolutional layers, where weights are determined by the gradient of confidence score for any target class with respect to the feature map entries. Attention rollout [1] measures it based on the attention weight of encoders of a Vision Transformer.
@@
Several recent studies revealed that a game-theoretic concept, Shapley values [24], is a powerful indicator of pixel contribution [18, 16, 8]. Shapley values were developed for equitable distribution of total rewards among players cooperating in game theory and quantify each player’s contribution by calculating the average difference in rewards with their presence versus absence. When applied to an image classifier, the pixels of an image are the players, which work cooperatively for the model output (e.g., confidence score). Unlike Grad-CAM and Attention rollout, Shapley values compute the contribution of pixels to the model output more directly. The former methods use feature maps or attention weights, the magnitude of whose entries are not necessarily well-aligned with the model output, whereas the latter uses logits or confidence scores. Indeed, Fig. 1 shows that the pixels with high Shapely values have a significantly larger impact on confidence scores than those determined by Grad-CAM or Attention rollout in both (a) insertion case and (b) deletion case. A crucial caveat of the aforementioned methods is that they identify a group of important pixels by the individual contribution of each pixel and overlook the collective contribution of multiple pixels. For example, Fig. 1(a) shows that the three methods only highlight the class object (i.e., duck) and do not indicate the background (i.e., sea) as an informative factor. However, the set of pixels with the highest contributions (e.g., highest Shapley values) does not imply the most informative pixel set as a whole because the information overlap among pixels is not considered. Indeed, the bottom row of Fig. 1(a) shows that the class object and background greatly impact in synergy the confidence score.
@@
Figure 1: Examples of image patches with high contributions to the output of ViT-T. (a) Starting from an empty image, image patches are inserted according to their contribution measured by each method. The numbers show the confidence score given the true class. (b) Starting from an original image, image patches are removed according to their contribution measured by each method. The numbers show the confidence score given the true class. The heatmaps highlight the image patches inserted (deleted) patches to obtain the correct (incorrect) classification. The patches selected are colored according to the timing of inclusion/deletion. For insertion, only the proposed method selects a patch from the background. For deletion, the proposed method highlights the class object only. For both cases, the proposed method highlights the least number of patches in the heatmaps. It sharply changes the confidence score, while others do not. As Grad-CAM is specialized for CNNs, Attention rollout performs better with the ViT architecture.
@@
In this paper, we propose an efficient game-theoretic visualization method of image pixels with a high impact on the prediction of an image classifier. Besides Shapley values, we exploit interactions, a game-theoretical concept that reflects the average effect of the cooperation of pixels. Unlike prior methods, including Grad-CAM, Attention rollout, and Shapley values, the proposed method takes into account the cooperative contribution of pixels and identifies the image pixels as a whole. Consequently, in Fig. 1(a), the proposed method identifies a pixel set on which the classifier puts high classification confidence. Similarly, in Fig. 1(b), it identifies a minimal pixel set without which the classification fails. Notably, we define self-context variants of Shapley values and interactions, and reduce the exponential times forwarding to linear times, which resolves the fundamental challenge of game-theoretic approaches to be handy tools for model explanation.
@@
In the experiments, we consider the insertion curve and deletion curve on a subset of ImageNet images that are correctly classified by a pretrained classifier. Starting from fully masked images, an insertion curve plots the increase of classification accuracy as unmasking image patches from highly contribution ones determined by each method. Similarly, a deletion curve plots the accuracy decrease from the clean images to fully masked ones. The results show that the proposed method gives sharp insertion/deletion curves. For example, the classification accuracy reached $90\%$ with images with 4% unmasked patches if selected by the proposed method, significantly outperforming the results of Grad-CAM (accuracy of $2\%$), Attention rollout (accuracy of $4\%$), and Shapley values (accuracy of $25\%$). Similar results are observed for the deletion curves and also when we use common corruptions instead of masking. Qualitatively, the heatmaps using the patches selected in the early stage of the insertion curve show that the proposed method using the insertion strategy highlights both a class object and background, while the other methods mostly highlight the class object only. Meanwhile, in the heatmaps from the deletion curves, the proposed method particularly highlights the class-discriminative region of the object, while the others do not. Our contributions are summarized as follows:
@@
  • Our analysis supports a simple greedy strategy from a game-theoretic perspective. This suggests the use of self-context variants of Shapley values and interactions, which can be computed exponentially faster than computing the original ones.
  • Extensive experiments show that our method more accurately identifies the pixels that are highly contributing to the model outputs than widely used visualization methods.
@@ 2 Related Work
Visualization of important image pixels for model explanation.
Various methods have been proposed to understand deep learning models for vision tasks by quantifying and visualizing the contribution of image pixels to the model output [31, 23, 1, 20, 18, 3, 5, 27, 6]. The contribution of pixels has been typically measured using feature maps in models.
For example, Grad-CAM [23] determines the contribution by applying weights to the feature maps of the convolutional layers of a CNN using gradients. Attention rollout [1], commonly used for Vision Transformers, calculates the contributions using attention maps. Several methods instead calculate the contribution of each pixel by analyzing the sensitivity of the confidence score with respect to each pixel [20, 18, 16, 8]. For example, RISE [20] (Randomized Input Sampling for Explanation) calculates the contributions empirically by probing the model with randomly masked images of the input image and obtaining the corresponding confidence scores. SHAP [18] (SHapley Additive exPlanations) distributes confidence scores fairly to contributions by leveraging Shapley values from game theory. Importantly, the aforementioned methods all measure the contribution of each pixel independently; the collection of important pixels consists of the pixels with high contributions. In contrast, this study identifies the important pixels by further taking into account the collective contributions of pixels.
@@
Several recent studies have utilized a game-theoretic concept, interactions, to analyze various phenomena of deep learning models and quantify an effect of pixel cooperation on the model inference [7, 9, 21, 28, 29, 25]. Wang et al. [28] showed that the transferability of adversarial images has a negative correlation to the interactions. Zhang et al. [29] showed the similarity between the computation of interactions and dropout regularization. Deng et al. [9] discussed the difference in information obtained between humans and machine learning models using interactions. Sumiyasu et al. [25] investigated misclassification by models using interactions and discovered that the distribution of interactions varies with the type of misclassified images. Thus, interactions are helpful for understanding the model from the perspective of cooperative relationships between pixels. A critical issue of interaction-based analysis is its computational cost; the computation of interaction requires an exponential number of forward passes with respect to the number of pixels. In this paper, we propose an efficient approach to explain a model using variants of interactions (and also Shapley values), achieving the identification of an important group of pixels with only a linear number of forward passes.
@@ 3 Preliminaries
Shapley values were proposed in game theory to measure the contribution of each player to the total reward that is obtained from multiple players working cooperatively [24]. Let $N=\{1,\ldots,n\}$ be the index set of players, and let $2^{N}\stackrel{{\scriptstyle\rm{def}}}{{=}}\{S\,|\,S\subseteq N\}$ be its power set. Given a reward function $f:2^{N}\to\mathbb{R}$, the Shapley value $\phi(i\,|\,N)$ of player $i$ with a context $N$ is defined as follows.
@@
In the application of Shapley values and interactions to image classifiers, an image $x$ with $n$ pixels is regarded as the index set $N=\{1,\ldots,m\}$ of players. Typically, the reward function $f$ is defined by $f(x)=\log\frac{P(y\,|\,x)}{1-P(y\,|\,x)}$ [9], where $y$ represent the class of $x$, and $P(y\,|\,x)$ denotes the classifier’s confidence score on class $y$ with input $x$. The reward $f(S)$ of a subset of pixels $S\subset 2^{N}$ of image $x$ is similarly computed by feeding a partially masked $x$ to the classifier (pixels in $N\setminus S$ are masked). If the classifier is a convolutional neural network (CNN), the masked region is conventionally filled with some base value, such as 0 and the average pixel value [2, 30]. Note that such a replacement may drop the original information of an image but also inject a new feature. Thus, the choice of base value affects the Shapley values and interactions. In contrast, when a Vision Transformer is used, one can realize masking in a rigid manner by applying a mask to the attention. To our knowledge, most prior studies exploited Shapley values and interactions on CNNs with the base value replacement, which might not unleash the full potential of these quantities. To our knowledge, the only exception is [8], which demonstrated that Shapley values can be calculated more accurately using attention masking. We follow this strategy in the computation of Shapley values and interactions for Vision Transformers.
@@ 4.1 Pixel insertion
By its formulation, this problem is an NP-hard problem in general. Particularly, $f$ is here a CNN or Vision Transformer, [^1] With this assumption, we use a slight abuse of notation and assume, e.g., $f(\{a,\{b,c\}\})=f(\{a,b,c\})$ because in either case of $\{a,\{b,c\}\}$ or $\{a,b,c\}$, we input the image with pixels $a,b,c$ to the model. a highly nonlinear function. Thus, we resort to a greedy strategy to solve it approximately. For $k=1$, the index $b_{1}\in N$ of the pixel with the highest Shapley value of $\phi(b_{1}\,|\,\{b_{1}\})$ gives the optimal set $S_{1}=\{b_{1}\}$ by the its definition.
For $k=2$, we select the next pixel $b_{2}$ with the one maximizing $f(\{b_{1},b_{2}\})$. Importantly, this is equivalent to maximizing the sum of the Shapley value and interaction, not the Shapley value alone.
@@
We refer to such a particular form of Shapley values and interactions to be self-context in the pixel insertion approach, and they play an essential role in our framework.
For $k\geq 3$, we can similarly show that maximizing $f(S_{k-1}\cup\{b\})$ with respect to $b_{k}$ is equivalent to
@@
Equation (7) shows that for the identification of index $b_{k}$ for $S_{k}$, it is essential to consider the interaction between $S_{k-1}$ and $b_{k}$. Even when a pixel indexed $b$ has a large Shapley value (the first term), it may have a large negative interaction (the second term) if its pixel information overlaps with that of $S_{k-1}$. Namely, collecting pixels with large Shapley values does not give the most informative pixel set.
To summarize, our analysis justifies a very simple greedy algorithm Algorithm 1 from a game-theoretic perspective. The algorithm seems trivial in hindsight, but prior studies visualize highly contribution pixels only using Shapley values [18, 16, 8].
@@
The identification of important pixels (or patches, in practice) using Shapley values requires $\mathcal{O}(|N|2^{|N|})$ times of forward passes because of the average over all $S\in N\setminus\{i\}$ for all $i\in N$ (cf. Eq. (1)). In contrast, our approach only requires $\mathcal{O}(|N|)$ times of forward passes.
@@
We now give an intuitive example for showing the necessity of interactions using Set-Sum task. Set-Sum task is a variant of Problem 1 with a collection of integers $N\subset\mathbb{Z}$ and reward function $f(S)=s$ for $S\subseteq N$, where $s$ denotes the sum of all types of integers in $S$. For example, $s=3$ for $S=\{2,2,1\}$. Note that for any $i\in N$, we have $f(S_{k-1}\cup\{i\})=S_{k-1}+i$ if $i\notin S_{k-1}$ and otherwise $f(S_{k-1}\cup\{i\})=S_{k-1}$. In this way, when the features already possessed are equal to the newly added features, the model does not gain new information. This shows the role of interaction in considering information redundancy.
@@
Through empirical evaluation, we confirm the advantage of using interactions in the visual Set-Sum task on the synthetic MNIST dataset. This task is to accurately predict the sum of all types of numbers in an image using a model. We constructed composite images, each of which consists of four randomly selected MNIST images arranged in a 2x2 grid (cf. Fig. 2(a)). The label of a composite image is the sum of all types of numbers in the image as in the Set-Sum problem. The evaluation metric utilizes the insertion curve, as detailed in Sec. 5. For the model and dataset details, refer to Appendix A. The insertion curves in Fig. 2(b) show that the MoXI achieves higher accuracy than the methods using MoXI(-), which uses self-context Shapley values, and the Shapley value methods when $50\%$ and $75\%$ of the image area are unmasked, i.e., the second and the third number is appended. Therefore, we demonstrate that MoXI acquires non-redundant information more effectively than the other methods.
Figure 2: (a) Example of the synthetic MNIST dataset in the visual Set-Sum task. The label of a composite image is the sum of all types of numbers in the image. (b) Insertion curves. The curves illustrate the change of accuracy when adding image patches gradually with high contributions identified by different methods at various unmasked image rates, ranging from 00 to $100\%$. These curves use a masking method that fills in zeros for game-theoretic calculations and model input during classification accuracy measurement. MoXI(-) employs self-context Shapley values, whereas MoXI combines these with interactions among highly-contributive patches.

### v2

@@ Abstract
To better understand the behavior of image classifiers, it is useful to visualize the contribution of individual pixels to the model prediction. In this study, we propose a method, MoXI (Model eXplanation by Interactions), that efficiently and accurately identifies a group of pixels with high prediction confidence. The proposed method employs game-theoretic concepts, Shapley values and interactions, taking into account the effects of individual pixels and the cooperative influence of pixels on model confidence. Theoretical analysis and experiments demonstrate that our method better identifies the pixels that are highly contributing to the model outputs than widely-used visualization by Grad-CAM, Attention rollout, and Shapley value. While prior studies have suffered from the exponential computational cost in the computation of Shapley value and interactions, we show that this can be reduced to quadratic cost for our task. The code is available at https://github.com/KosukeSumiyasu/MoXI.
@@ 1 Introduction
Visualization of important image pixels has been widely used to understand machine learning models in computer vision tasks such as image classification [31, 23, 1, 20, 18, 3]. To this end, visualization methods compute the contribution of each pixel to model decisions. For example, Grad-CAM [23] measures the contribution using a weighted sum of the feature maps of convolutional layers, where weights are determined by the gradient of confidence score for any target class with respect to the feature map entries. Attention rollout [1] measures it based on the attention weight of encoders of a Vision Transformer.
@@
Several recent studies revealed that a game-theoretic concept, Shapley values [24], is a powerful indicator of pixel contribution [18, 16, 8]. In multi-player games, Shapley values measures the contribution of each player from the average change in the total game reward with his/her presence versus absence. When applied to an image classifier, the pixels of an image are the players, which work cooperatively for the model output (e.g., confidence score). Unlike Grad-CAM and Attention rollout, Shapley values compute the contribution of pixels to the model output more directly. The former methods use feature maps or attention weights, the magnitude of whose entries are not necessarily well-aligned with their contributions to the model output, whereas the latter uses logits or confidence scores. Indeed, Fig. 1 shows that the pixels with high Shapely values have a significantly larger impact on confidence scores than those determined by Grad-CAM or Attention rollout in both (a) insertion case and (b) deletion case.
A crucial caveat of the aforementioned methods is that they identify a group of important pixels by the individual contribution of each pixel and overlook the collective contribution of multiple pixels. For example, Fig. 1(a) shows that the three methods only highlight the class object (i.e., duck) and do not indicate the background (i.e., sea) as an informative factor. However, the set of pixels with the highest contributions (e.g., highest Shapley values) does not imply the most informative pixel set as a whole because the information overlap among pixels is not considered. Indeed, the bottom row of Fig. 1(a) shows that the class object and background greatly impact in synergy the confidence score.
@@
Figure 1: Examples of image patches with high contributions to the output of ViT-T. (a) Starting from an empty image, image patches are inserted according to their contribution measured by each method. (b) Starting from an original image, image patches are removed according to their contribution measured by each method. The heatmaps highlight the image patches inserted (deleted) to obtain the correct (incorrect) classification. The selected patches are colored according to the timing of insertion/deletion. For insertion, only the proposed method selects patches from the background. For deletion, the proposed method highlights the class object only. For both cases, the proposed method highlights the least number of patches while achieving the highest/lowest confidence score.
@@
In this paper, we propose an efficient game-theoretic visualization method of image pixels with a high impact on the prediction of an image classifier. Besides Shapley values, we exploit interactions, a game-theoretical concept that reflects the average effect of the cooperation of pixels. Namely, unlike prior methods, including Grad-CAM, Attention rollout, and Shapley values, the proposed method takes into account the cooperative contribution of pixels and identifies the image pixels as a whole. In Fig. 1(a), the proposed method identifies a pixel set on which the classifier puts high classification confidence. Similarly, in Fig. 1(b), it identifies a minimal pixel set without which the classification fails. Notably, we define self-context variants of Shapley values and interactions, and reduce the number of forward passes from exponential to quadratic times, which resolves the fundamental challenge of game-theoretic approaches to be handy tools for model explanation.
@@
In the experiments, we consider the insertion curve and deletion curve on a subset of ImageNet images that are correctly classified by a pretrained classifier. Starting from fully masked images, an insertion curve plots the increase of classification accuracy as unmasking image patches from highly contributing ones determined by each method. Similarly, a deletion curve plots the accuracy decrease from the clean images to fully masked ones. The results show that the proposed method gives sharp insertion/deletion curves. For example, the classification accuracy reached $90\%$ with images with 4% unmasked patches if selected by the proposed method, significantly outperforming the results of Grad-CAM (accuracy of $2\%$), Attention rollout (accuracy of $4\%$), and Shapley values (accuracy of $25\%$). Similar results are observed for the deletion curves and also when we use common corruptions [15] instead of masking. Qualitatively, the heatmaps using the patches selected in the early stage of the insertion curve show that our method highlights both a class object and background, while the other methods mostly highlight the class object only. Meanwhile, in the heatmaps from the deletion curves, our method particularly highlights the class-discriminative region of the object, while the others do not.
Our contributions are summarized as follows:
@@
  • Our analysis supports a simple greedy strategy from a game-theoretic perspective, leading us to use self-context variants of Shapley values and interactions, which can be computed exponentially faster than computing the original ones.
  • Extensive experiments show that our method more accurately identifies the pixels that are highly contributing to the model outputs than standard visualization methods.
@@ 2 Related Work
Visual explanation of model decision.
Various methods have been proposed to understand deep learning models for vision tasks by quantifying and visualizing the contribution of image pixels to the model output [31, 23, 1, 20, 18, 3, 5, 27, 6]. The contribution of pixels has been typically measured using feature maps in models. For example, Grad-CAM [23] determines the contribution by applying weights to the feature maps of the convolutional layers of a CNN using gradients. Attention rollout [1], commonly used for Vision Transformers, calculates the contributions using attention maps. Several methods instead calculate the contribution of each pixel by analyzing the sensitivity of the confidence score with respect to each pixel [20, 18, 16, 8]. For example, RISE (Randomized Input Sampling for Explanation; [20]) calculates the contributions empirically by probing the model with randomly masked images of the input image and obtaining the corresponding confidence scores. SHAP (SHapley Additive exPlanations; [18]) distributes confidence scores fairly to contributions by leveraging Shapley values from game theory. Importantly, the aforementioned methods all measure the contribution of each pixel independently; the collection of important pixels consists of the pixels with high contributions. In contrast, this study identifies the important pixels by further taking into account the collective contributions of pixels.
@@
Several recent studies have utilized a game-theoretic concept, interactions, to analyze various phenomena of deep learning models and quantify an effect of pixel cooperation on the model inference [7, 9, 21, 28, 29, 25]. Wang et al. [28] showed that the transferability of adversarial images has a negative correlation to the interactions. Zhang et al. [29] showed the similarity between the computation of interactions and dropout regularization. Deng et al. [9] discussed the difference in information obtained between humans and machine learning models using interactions. Sumiyasu et al. [25] investigated misclassification by models using interactions and discovered that the distribution of interactions varies with the type of misclassified images. Thus, interactions are helpful for understanding the model from the perspective of cooperative relationships between pixels. A critical issue of interaction-based analysis is its computational cost; the computation of interaction requires an exponential number of forward passes with respect to the number of pixels. In this paper, we propose an efficient approach to explain a model using variants of interactions (and also Shapley values), achieving the identification of important pixels with only a quadratic number of forward passes.
@@ 3 Preliminaries
Shapley values was proposed in game theory to measure the contribution of each player to the total reward that is obtained from multiple players working cooperatively [24]. Let $N=\{1,\ldots,n\}$ be the index set of players, and let $2^{N}\stackrel{{\scriptstyle\rm{def}}}{{=}}\{S\,|\,S\subseteq N\}$ be its power set. Given a reward function $f:2^{N}\to\mathbb{R}$, the Shapley value $\phi(i\,|\,N)$ of player $i$ with a context $N$ is defined as follows.
@@
In the application of Shapley values and interactions to image classifiers, an image $x$ with $n$ pixels is regarded as the index set $N=\{1,\ldots,m\}$ of players. Typically, the reward function $f$ is defined by $f(x)=\log\frac{P(y\,|\,x)}{1-P(y\,|\,x)}$ [9], where $y$ represents the class of $x$, and $P(y\,|\,x)$ denotes the classifier’s confidence score on class $y$ with input $x$. The reward $f(S)$ of a subset of pixels $S\subset 2^{N}$ of image $x$ is similarly computed by feeding a partially masked $x$ to the classifier (i.e., the pixels in $N\setminus S$ are masked).
If the classifier is a convolutional neural network (CNN), the masked region is conventionally filled with some base value, such as 0 or the average pixel value [2, 30]. Such a replacement may drop the original information of an image but also inject a new feature. Thus, the choice of base value affects the Shapley values and interactions. In contrast, when a Vision Transformer is used, one can realize masking in a rigid manner by applying a mask to the attention. To our knowledge, most prior studies exploited Shapley values and interactions on CNNs with the base value replacement, which might not unleash the full potential of these quantities. To our knowledge, the only exception is [8], which demonstrated that Shapley values can be calculated more accurately using attention masking. We follow this strategy in the computation of Shapley values and interactions for Vision Transformers.
@@ 4.1 Pixel Insertion
By its formulation, this problem is an NP-hard problem in general. Particularly, $f$ is here a CNN or Vision Transformer, [^1] With this assumption, we use a slight abuse of notation and assume, e.g., $f(\{a,\{b,c\}\})=f(\{a,b,c\})$ because in either case of $\{a,\{b,c\}\}$ or $\{a,b,c\}$, we input the image with pixels $a,b,c$ to the model. a highly nonlinear function. Thus, we resort to a greedy strategy to solve it approximately.
For $k=1$, the index $b_{1}\in N$ of the pixel with the highest Shapley value of $\phi(b_{1}\,|\,\{b_{1}\})$ gives the optimal set $S_{1}=\{b_{1}\}$ by the its definition. For $k=2$, we select the next pixel $b_{2}$ with the one maximizing $f(\{b_{1},b_{2}\})$. Importantly, this is equivalent to maximizing the sum of the Shapley value and interaction, not the Shapley value alone.
@@
We refer to such a particular form of Shapley values and interactions to be self-context in the pixel insertion approach, and they play an essential role in our framework. For $k\geq 3$, we can similarly show that maximizing $f(S_{k-1}\cup\{b_{k}\})$ with respect to $b_{k}$ is equivalent to
@@
Equation (7) shows that for identifying of index $b_{k}$ for $S_{k}$, it is crucial to consider the interaction between $S_{k-1}$ and $b_{k}$. Even when a pixel indexed $b$ has a large Shapley value (the first term), it may have a large negative interaction (the second term) if its pixel information overlaps with that of $S_{k-1}$. Namely, collecting pixels with large Shapley values does not necessarily give the most informative pixel set.
To summarize, our analysis justifies a very simple greedy algorithm Algorithm 1 from a game-theoretic perspective. The algorithm seems trivial in hindsight, but prior studies visualize highly contributing pixels only using Shapley values [18, 16, 8].
@@
The identification of important pixels (or patches, in practice) using Shapley values requires $\mathcal{O}(|N|2^{|N|})$ times of forward passes because of the average over all $S\in N\setminus\{i\}$ for all $i\in N$ (cf. Eq. (1)). In contrast, our approach only requires $\mathcal{O}(|N|^{2})$ times of forward passes in the worst case (see Appendix C for details of the algorithm complexity and runtime).
@@
We now give an intuitive example for showing the necessity of interactions using Set-Sum task. Set-Sum task is a variant of Problem 1 with a collection of integers $N\subset\mathbb{Z}$ and reward function $f(S)=s$ for $S\subseteq N$, where $s$ denotes the sum of all types of integers in $S$. For example, $s=3$ for $S=\{2,2,1\}$. Note that for any $i\in N$, we have $f(S_{k-1}\cup\{i\})=f(S_{k-1})+i$ if $i\notin S_{k-1}$ and otherwise $f(S_{k-1}\cup\{i\})=f(S_{k-1})$. In this way, when the features already possessed are equal to the newly added features, the model does not gain new information. This shows the role of interaction in considering information redundancy.
@@
We empirically confirm the advantage of using interactions in the visual Set-Sum task on the synthetic MNIST dataset. This task is to accurately predict the sum of all types of numbers in an image using a model. We constructed composite images, each of which consists of four randomly selected MNIST images arranged in a 2x2 grid (cf. Fig. 2(a)). The label of a composite image is the sum of all types of numbers in the image as in the Set-Sum problem. The evaluation metric utilizes the insertion curve, as detailed in Sec. 5. For the model and dataset details, refer to Appendix A. The insertion curves in Fig. 2(b) show that the MoXI achieves higher accuracy than the methods using MoXI(-), which uses self-context Shapley values, and the Shapley value methods when $50\%$ and $75\%$ of the image area are unmasked, i.e., the second and the third number is appended. This demonstrates that MoXI acquires non-redundant information more effectively.
Figure 2: (a) Example of a synthetic MNIST image in the visual Set-Sum task, labeled 17 by the sum of all types of numbers in the image. (b) Insertion curves. The curves illustrate the change of accuracy when adding image patches gradually with high contributions identified by different methods at various unmasked image rates, ranging from 00 to $100\%$. These curves use a masking method that fills in zeros for game-theoretic calculations and model input during classification accuracy measurement. MoXI(-) only employs self-context Shapley values, whereas MoXI additionally uses interactions across highly contributing patches.